### v2

# On the Effects of Modeling Errors on Distributed Continuous-time Filtering

Abstract
This paper offers a comprehensive performance analysis of the distributed continuous-time filtering in the presence of modeling errors. First, we introduce two performance indices, namely the nominal performance index and the estimation error covariance. By leveraging the nominal performance index and the Frobenius norm of the modeling deviations, we derive the bounds of the estimation error covariance and the lower bound of the nominal performance index. Specifically, we reveal the effect of the consensus parameter on both bounds. We demonstrate that, under specific conditions, an incorrect process noise covariance can lead to the divergence of the estimation error covariance. Moreover, we investigate the properties of the eigenvalues of the error dynamical matrix. Furthermore, we explore the magnitude relations between the nominal performance index and the estimation error covariance. Finally, we present some numerical simulations to validate the effectiveness of the theoretical results.
keywords:
Distributed continuous-time filtering; Modeling error; Performance analysis; Divergence analysis; Sensor networks.
Sensor networks have been extensively employed in various domains, including satellite navigation [10], environmental monitoring [11], and robot cooperative mapping [18]. Distributed state estimation [13, 14, 2, 3] plays a significant role in sensor networks. It is utilized to estimate the system state for each intelligent sensor within sensor networks by leveraging the collective information from its own observations as well as the shared knowledge from its local neighbors.
Distributed state estimation can be classified into two primary categories: distributed continuous-time filter and distributed discrete-time filter, which are based on two distinct types of state space model frameworks. In comparison to the distributed discrete-time filters, fewer research advancements have been achieved in the stochastic continuous-time model setting, primarily due to its inherent complexity. Some results of distributed continuous-time state estimation are presented as follows. Olfati-Saber [13] proposed a distributed continuous-time Kalman filter by integrating the consensus fusion term of the state estimate into the Kalman-Bucy filter [7]. It is important to note that a crucial requirement for this distributed filter was the availability of the prior knowledge of the initial state and covariance, and its stability was proven under the assumption of a noise-free scenario. To tackle the issue of unguaranteed boundedness of covariance matrices in the absence of each sensor’s observability[13], Kim et al. [8] proposed algorithms by exchanging covariance matrices among sensors. However, the stability and the parameter conditions were also derived under the noise-free assumption. Similarly to the filter structure presented in [13], Wu et al. [20] utilized the inverse of covariance matrices as weights for the consensus terms. Nonetheless, this method encountered similar issues as discussed above. Battilotti et al. [1] exhibited an asymptotically optimal distributed filter, and provided a comprehensive analysis of its convergence and optimality. In this paper, the fundamental filter framework for performance analysis in the presence of modeling errors will be based on the distributed continuous-time filter presented in [1]. This choice is motivated by the less restrictive stability conditions and the asymptotically optimal properties exhibited by this filter. It is worth mentioning that the analysis methods employed in this framework can be adapted to other filters with ease.
In practical scenarios, obtaining precise models is rarely feasible, and modeling errors are widespread. The modeling errors involve deviations in the state matrix, deviations in the measurement matrix, and mismatched noise covariances. Such modeling errors have the potential to result in a decline in the filter performance and even lead to system failure. Several studies have explored the performance of continuous-time filters for a single sensor when confronted with modeling errors. T. Nishimura [12] established a conservative design criterion to ensure the upper bound of the estimation error covariance for continuous-time systems with incorrect noise covariances. Fitzgerald [4] investigated into the divergence of the Kalman filter, highlighting the mean square error may become unbounded due to the incorrect process noise covariance. Toda and Patel [19] derived performance index bounds and mean square error bounds for the suboptimal filters, considering the presence of modeling errors. Sangsuk-Iam and Bullock [15] analyzed the behavior of the continuous-time Kalman filter under mismatched noise covariances. Additionally, some existing literature on state estimation addresses arbitrary noise scenarios [17, 5, 9]. However, these studies primarily focus on the single-sensor systems. When shifting the focus to the distributed filters, numerous challenges arise. These challenges include dealing with the new distributed filter structure, addressing the coupling terms, and handling the behavior of the consensus parameter. This paper will investigate the effect of the modeling errors on the distributed continuous-time filters.
Motivated by the aforementioned observations, the primary objective of this paper is to offer a comprehensive performance analysis of the distributed continuous-time filter in the presence of modeling errors. This analysis aims to enrich individuals’ comprehension and prognostication of the filters’ behavior. The main contributions of this paper are summarized as follows:
  1. 1.
    We define two performance indices, namely the nominal performance index and the estimation error covariance, as metrics to assess the performance of the filter. By utilizing the nominal performance index and the norm of the modeling deviations, the bounds of the estimation error covariance are determined (Theorem 1). In particular, the asymptotic decay behavior of this bound in terms of $\gamma_{u}$ is revealed. These results play an important role in evaluating the the estimation error covariance by utilizing the nominal models in the presence of the modeling errors.
  2. 2.
    We reveal that an incorrect process noise covariance can lead to the divergence of the estimation error covariance for undirected connected networks, regardless of the magnitude of the consensus parameter (Theorem 2). It is demonstrated that the nominal parameters must be reasonably designed to ensure the convergence of the estimation error covariance.
  3. 3.
    The relations between the nominal performance index and the estimation error covariance, including their magnitude relations and the norm bound of their difference, are analyzed based on the characteristics of the noise covariance deviations (Theorem 3). This result offers insights into the judicious selection of the nominal covariance to ensure a conservative estimation.
The subsequent sections of this paper are structured as follows. Section 1 provides the problem statement. Section 2 presents the nominal distributed filter and the estimation error covariance. Section 3 focuses on analyzing bounds of the estimation error covariances, divergence, and relations between different performance indices. Section 4 exhibits some examples to validate the effectiveness of the theoretical results. Section 5 concludes this paper.
[…]
1 Problem Statement
Consider a continuous-time linear stochastic system, measured by a sensor network of $N$ sensors, described by
$\dot{x}(t)=Ax(t)+\omega(t),$ (1)
$y_{i}(t)=C_{i}x(t)+\nu_{i}(t),~{}~{}~{}i=1,2,...,N,$
where $t$ is the time index, $x(t)\in\mathcal{R}^{n}$ is the system state, $y_{i}(t)\in\mathcal{R}^{m_{i}}$ is the measurement of sensor $i$, $A\in\mathcal{R}^{n\times n}$ is the state matrix, $C_{i}\in\mathcal{R}^{m_{i}\times n}$ is the measurement matrix of sensor $i$, $\omega(t)\in\mathcal{R}^{n}$ is the process noise, and $\nu_{i}(t)\in\mathcal{R}^{m_{i}}$ is the measurement noise of sensor $i$. It is assumed that $\omega(t)$ and $\nu_{i}(t)$ are zero-mean white noise processes, and these noise processes are uncorrelated with each other and with the initial state $x(0)$, i.e., $\mathbb{E}\{\omega(t)\omega^{T}(\tau)\}=Q\delta(t-\tau)$, $\mathbb{E}\{\nu_{i}(t)\nu^{T}_{i}(\tau)\}=R_{i}\delta(t-\tau),~{}i=1,\ldots,N,$ $\mathbb{E}\{\omega(t)\nu^{T}_{i}(\tau)\}=0,~{}i=1,\ldots,N,$ and $\mathbb{E}\{\nu_{i}(t)\nu^{T}_{j}(\tau)\}=0,~{}i,j=1,\ldots,N,i\neq j$, $Q\in\mathcal{R}^{n\times n}$ and $R_{i}\in\mathcal{R}^{m_{i}\times m_{i}}$ denote the process noise covariance and the measurement noise covariance of sensor $i$, respectively, and $\delta(t)$ represents the Dirac delta function, given by $\delta(t-\tau)=1$ if $t=\tau$ and 0 otherwise. The communication topology of sensor networks is described by the graph theory in Appendix A.
We adopt a classical distributed continuous-time Kalman filter in [1], given by
$\dot{\hat{x}}_{i}(t)=$ $A\hat{x}_{i}(t)+K_{i}(y_{i}(t)-C_{i}\hat{x}_{i}(t))$ (2)
$+\gamma P(\infty)\sum_{j\in\mathcal{N}_{i}}(\hat{x}_{j}(t)-\hat{x}_{i}(t)),$
where $K_{i}=NP(\infty)C^{T}_{i}R^{-1}_{i}$, $P(\infty)$ is the solution of
$0=$ $AP(\infty)+P(\infty)A^{T}+Q-P(\infty)C^{T}_{c}R^{-1}_{d}C_{c}P(\infty),$ (3)
$C_{c}=[C^{T}_{1},\ldots,C^{T}_{N}]^{T}$ is the augmented measurement matrix, and $R_{d}=\text{diag}\{R_{1},\ldots,R_{N}\}$ is the augmented measurement noise covariance.
Accurate model parameters, such as $A$, $C_{i}$, $Q$, and $R_{i}$, are often difficult to obtain in practice. Therefore, we utilize the nominal model parameters, denoted as $A_{u}$, $C_{i,u}$, $Q_{u}$, and $R_{i,u}$, in the practical filter. The deviations between the nominal parameters and the actual parameters are defined as
$\Delta A=A_{u}-A,~{}~{}~{}\Delta C_{i}=C_{i,u}-C_{i},$ (4)
$\Delta Q=Q_{u}-Q,~{}~{}~{}\Delta R_{i}=R_{i,u}-R_{i}.~{}~{}$
For these deviations, we may have some prior knowledge or estimates, such as the Frobenius norm bound of these deviations and the magnitude relations between the nominal and actual parameters. The objective of this paper is to analyze the effect of these modeling errors on the performance of the distributed filter and provide performance estimates based on the available prior information. The problem of the effects of the modeling errors on distributed continuous-time filtering is formulated as follows:
  1. 1.
    Based on the prior Frobenius norm bound for the deviations between the nominal and actual parameters, we evaluate the performance of the distributed filter by utilizing the known nominal model parameters.
  2. 2.
    Explain how an incorrect process noise covariance can lead to the divergence of the distributed filter, even when the system model is accurate. Investigate the relations between different performance indices in terms of the consensus gain.
2 Performance Indices
This section provides two performance indices, namely the nominal performance index and the estimation error covariance, and the estimation error covariance is derived based on the nominal distributed Kalman filter.
2.1 Nominal Distributed Kalman Filter
In practice, we utilize the nominal parameters $A_{u}$, $C_{i,u}$, $Q_{u}$, and $R_{i,u}$ in the distributed filter. Similarly to the ideal distributed Kalman filter (2), the nominal distributed Kalman filter with known and inaccurate nominal parameters can be expressed as
$~{}~{}~{}~{}~{}~{}\dot{\hat{x}}_{i,u}(t)=$ $A_{u}\hat{x}_{i,u}(t)+K_{i,u}(y_{i}(t)-C_{i,u}\hat{x}_{i,u}(t))$ (5)
$+\gamma_{u}P_{u}(\infty)\sum_{j\in\mathcal{N}_{i}}(\hat{x}_{j,u}(t)-\hat{x}_{i,u}(t)),$
where $K_{i,u}=NP_{u}(\infty)C^{T}_{i,u}R^{-1}_{i,u}$, $P_{u}(\infty)$ is the solution of
$0=$ $A_{u}P_{u}(\infty)+P_{u}(\infty)A^{T}_{u}+Q_{u}$ (6)
$-P_{u}(\infty)C^{T}_{c,u}R^{-1}_{d,u}C_{c,u}P_{u}(\infty),$
[…]
The compact form of (5) can be written as
$\dot{\hat{x}}_{c,u}(t)=\mathcal{A}_{u}\hat{x}_{c,u}(t)+K_{d,u}y_{c}(t),$ (7)
where $G_{i,u}=A_{u}-K_{i,u}C_{i,u}$, $G_{d,u}=\text{diag}\{G_{1,u},\ldots,G_{N,u}\}$, $K_{d,u}=\text{diag}\{K_{1,u},\ldots,K_{N,u}\}$, and
$\mathcal{A}_{u}=G_{d,u}-\gamma_{u}(\mathcal{L}\otimes P_{u}(\infty)).{}{}$ (8)
For the nominal distributed filter, we define $\Sigma_{u}(t)$ as the nominal performance index, which evaluates the filter’s performance based on the nominal parameters:
$~{}~{}\dot{\Sigma}_{u}(t)=$ $\mathcal{A}_{u}\Sigma_{u}(t)+\Sigma_{u}(t)\mathcal{A}^{T}_{u}$ (9)
$+K_{d,u}R_{d,u}K^{T}_{d,u}+U_{N}\otimes Q_{u},$
where $U_{N}=1_{N}1^{T}_{N}$. The nominal performance index is derived under the assumption that there are no deviations in the nominal parameters. It has the same form as the estimation error covariance presented in [1].
Remark 1.
In practical engineering applications, the consensus gain can be directly adjusted in the code. As the consensus gain increases, indicating that more information is being utilized from other filters, the ideal distributed filter is expected to achieve better performance. However, due to modeling errors, the effect of the consensus gain requires further exploration in the following.
2.2 Estimation Error Covariance
This subsection derives the bounds of the estimation error covariance by utilizing the nominal performance index and the norm of the modeling deviations, and presents the lower bound of the nominal performance index. Particularly, the effect of the consensus parameter on both bounds is revealed.
[…]
Proposition 1.
Consider the system (1) and the nominal distributed filter (5). The estimation error, the augmented estimation error, and the augmented estimation error covariance can be expressed as follows:
  1. 1.
    The estimation error is
    $~{}~{}~{}~{}\dot{\eta}_{i}(t)$ $=F_{i,u}x(t)+G_{i,u}\eta_{i}(t)+\omega(t)-K_{i,u}\nu_{i}(t)$ (10)
    $~{}~{}~{}~{}-\gamma_{u}P_{u}(\infty)\sum_{j\in\mathcal{N}_{i}}(\eta_{i}(t)-\eta_{j}(t)),$
    where $F_{i,u}=A-A_{u}-K_{i,u}(C_{i}-C_{i,u})$.
  2. 2.
    The augmented estimation error is
    $\dot{\eta}_{c}(t)$ $=F_{d,u}(1_{N}\otimes x(t))+\mathcal{A}_{u}\eta_{c}(t)$ (11)
    $~{}~{}~{}~{}+1_{N}\otimes\omega(t)-K_{d,u}\nu_{c}(t),~{}~{}~{}~{}~{}$
    where $F_{d,u}=\text{diag}\{F_{1,u},\ldots,F_{N,u}\}$.
  3. 3.
    The augmented estimation error covariance is
    $\dot{\Sigma}_{e}(t)$ $=\mathcal{A}_{u}\Sigma_{e}(t)+\Sigma_{e}(t)\mathcal{A}^{T}_{u}+F_{d,u}S^{T}(t)$ (12)
    $~{}~{}~{}~{}+S(t)F^{T}_{d,u}+K_{d,u}R_{d}K^{T}_{d,u}+U_{N}\otimes Q,$
    where $U_{N}=1_{N}1^{T}_{N}$, $\dot{S}(t)=\mathcal{A}_{u}S(t)+S(t)A^{T}_{d}+F_{d,u}X(t)+U_{N}\otimes Q,$ and $\dot{X}(t)=A_{d}X(t)+X(t)A^{T}_{d}+U_{N}\otimes Q$.
Proof.
Item 1) and Item 2): For computational simplicity, the time index is omitted. By combining (1) with (5), one has
$\dot{\eta}_{i}$ $=\dot{x}-\dot{\hat{x}}_{i,u}$ (13)
$=(Ax+\omega)-\Big{(}A_{u}\hat{x}_{i,u}+K_{i,u}(y_{i}-C_{i,u}\hat{x}_{i,u})$
$~{}~{}~{}~{}+\gamma_{u}P_{u}(\infty)\sum_{j\in\mathcal{N}_{i}}(\hat{x}_{j,u}-\hat{x}_{i,u})\Big{)}$
$=(A-A_{u})x+A_{u}(x-\hat{x}_{i,u})+\omega$
$~{}~{}~{}~{}-K_{i,u}C_{i,u}(x-\hat{x}_{i,u})-K_{i,u}(C_{i}-C_{i,u})x$
$~{}~{}~{}~{}-K_{i,u}\nu_{i}-\gamma_{u}P_{u}(\infty)\sum_{j\in\mathcal{N}_{i}}((x-\hat{x}_{i,u})-(x-\hat{x}_{j,u}))$
$=(A-A_{u})x+A_{u}\eta_{i}+\omega-K_{i,u}C_{i,u}\eta_{i}-K_{i,u}\nu_{i}$
$~{}~{}~{}~{}-K_{i,u}(C_{i}-C_{i,u})x-\gamma_{u}P_{u}(\infty)\sum_{j\in\mathcal{N}_{i}}(\eta_{i}-\eta_{j}).$
According to the definition of $F_{i,u}$ and $G_{i,u}$, (13) can be rewritten as (10). Next, using the definition of $F_{d,u}$, $G_{d,u}$, and $K_{d,u}$ above (8), the augmented estimation error can be expressed as:
$\dot{\eta}_{c}(t)$ $=F_{d,u}(1_{N}\otimes x(t))+G_{d,u}\eta_{c}(t)$
$~{}~{}~{}~{}+1_{N}\otimes\omega(t)-K_{d,u}\nu_{c}(t)-\gamma_{u}(\mathcal{L}\otimes P_{u}(\infty))\eta_{c}(t).$
Finally, by denoting $\mathcal{A}_{u}=G_{d,u}-\gamma_{u}(\mathcal{L}\otimes P_{u}(\infty))$, (11) can be obtained.
Item 3): Define $\xi=[\eta^{T}_{c},(1_{N}\otimes x)^{T}]^{T}$ [6], and it follows $\dot{\xi}=\mathcal{F}\xi+\mathcal{B}\rho,$ where $\rho=[\nu^{T}_{c},(1\otimes\omega)^{T}]^{T}$,
$\mathcal{F}=\left[\begin{array}[]{cc}\mathcal{A}_{u}&F_{d,u}\\ 0&A_{d}\end{array}\right],{}$ (14)
and
$\mathcal{B}=\left[\begin{array}[]{cc}-K_{d,u}&I\\ 0&I\end{array}\right].$ (15)
Define $\Sigma_{\xi}=\mathbb{E}\{\xi\xi^{T}\}$ and $\Phi=\mathbb{E}\{\rho\rho^{T}\}$, and one has
$\dot{\Sigma}_{\xi}=\mathcal{F}\Sigma_{\xi}+\Sigma_{\xi}\mathcal{F}^{T}+\mathcal{B}\Phi\mathcal{B}^{T},$ (16)
and
$\Phi=\left[\begin{array}[]{cc}R_{d}&0\\ 0&U_{N}\otimes Q\end{array}\right].{}{}{}{}{}$ (17)
Finally, the term $\dot{\Sigma}_{e}$ can be obtained based on (16), and the detailed expression is given in (12).
It is worth mentioning that if $F_{d,u}=0$, the estimation error covariance (12) is simplified as
$\dot{\Sigma}_{e}(t)$ $=\mathcal{A}_{u}\Sigma_{e}(t)+\Sigma_{e}(t)\mathcal{A}^{T}_{u}+K_{d,u}R_{d}K^{T}_{d,u}+U_{N}\otimes Q.$ (18)
3 Performance Analysis
This section provides an in-depth performance analysis for the distributed continuous-time filtering from multiple different aspects: bounds of the estimation error covariance, divergence analysis, and relations between the different performance indices. First, some assumptions are presented here for further analysis.
Assumption 1.
The communication graph is undirected and connected.
Assumption 2.
$(A_{u},C_{c,u})$ is observable, and $(A_{u},Q^{1/2}_{u})$ is controllable.
Proposition 2.
[1] Consider the system (1) and the nominal distributed filter (5). Suppose that Assumptions 1 and 2 hold. There exists $\bar{\gamma}_{u_{0}}$ such that for all $\gamma_{u}>\bar{\gamma}_{u_{0}}$, $\mathcal{A}_{u}$ is Hurwitz stable. The $\bar{\gamma}_{u_{0}}$ is given by
$\bar{\gamma}_{u_{0}}=\frac{\|P^{-1}_{u}(\infty)A_{u}+A^{T}_{u}P^{-1}_{u}(\infty)\|_{2}+\tilde{\gamma}_{u_{0}}}{\lambda_{N-1}(\mathcal{L})},$ (19)
where $\tilde{\gamma}_{u_{0}}=4N^{2}\alpha((P^{-1}_{u}(\infty)Q_{u}P^{-1}_{u}(\infty)+C^{T}_{c,u}R^{-1}_{d,u}\\ C_{c,u})^{-1})\alpha(C^{T}_{c,u}R^{-1}_{d,u}C_{c,u})$, $\alpha(\cdot)$ is the spectral abscissa as defined in the Notations, and $\gamma_{u}$ and $\mathcal{A}_{u}$ are defined in (5) and (8), respectively.
Given Proposition 2, select any $\gamma_{u_{0}}$ satisfying $\gamma_{u_{0}}>\bar{\gamma}_{u_{0}}$ as the initial consensus parameter. Define the parameter $\mathcal{A}_{u_{0}}$ as
$\mathcal{A}_{u_{0}}=G_{d,u}-\gamma_{u_{0}}(\mathcal{L}\otimes P_{u}(\infty)).$ (20)
It is worth mentioning that these assumptions are primarily provided for the nominal parameters rather than the actual parameters. Assumption 1 is the basic assumption of the communication networks. Assumption 2 provides the controllability and observability of the nominal model. Proposition 2 provides a lower bound for the consensus gain to ensure that the nominal matrix $\mathcal{A}_{u}$ is Hurwitz stable.
3.1 Bounds of the Estimation Error Covariance
This subsection provides the bounds of the estimation error covariance by utilizing the nominal performance index and the parameter deviations, and presents the lower bound of the nominal performance index. Moreover, it reveals the effect of the consensus parameter on both bounds.
Assumption 3.
If $F_{d,u}\neq 0$, $A$ is a Hurwitz matrix. If $F_{d,u}=0$, no additional conditions apply.
Remark 2.
For Assumption 3, if $F_{d,u}\neq 0$ and $A$ is unstable, the estimation error covariance will diverge, making it meaningless to investigate the corresponding result. Hence, Assumption 3 has been made to provide a unified expression and analysis.
Proposition 3.
Consider the system (1) and the nominal distributed filter (5). For $\Sigma_{u}(t)$ and $\Sigma_{e}(t)$, it holds
  1. 1.
    Under Assumptions 1-2, with $\gamma_{u}\geq\gamma_{u_{0}}$ defined in (19), $\Sigma_{u}(t)$ converges to $\bar{\Sigma}_{u}$, where $\bar{\Sigma}_{u}$ is the solution of the following Lyapunov equation
    $\mathcal{A}_{u}\bar{\Sigma}_{u}+\bar{\Sigma}_{u}\mathcal{A}^{T}_{u}+K_{d,u}R_{d,u}K^{T}_{d,u}+U_{N}\otimes Q_{u}=0.$ (21)
  2. 2.
    Under Assumptions 1-3, with $\gamma_{u}\geq\gamma_{u_{0}}$ defined in (19), $\Sigma_{e}(t)$ converges to $\bar{\Sigma}_{e}$, where $\bar{\Sigma}_{e}$ is the solution of the following Lyapunov equation
    $\mathcal{A}_{u}\bar{\Sigma}_{e}+\bar{\Sigma}_{e}\mathcal{A}^{T}_{u}+F_{d,u}\bar{S}^{T}+\bar{S}F^{T}_{d,u}$ (22)
    $~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}+K_{d,u}R_{d}K^{T}_{d,u}+U_{N}\otimes Q=0,$
    where $\bar{S}$ and $\bar{X}$ are the solutions of the following equations $\mathcal{A}_{u}\bar{S}+\bar{S}A^{T}_{d}+F_{d,u}\bar{X}+U_{N}\otimes Q=0$ and $A_{d}\bar{X}+\bar{X}A^{T}_{d}+U_{N}\otimes Q=0$, respectively.
Proof.
Item 1: Based on the results in Proposition 2, it can be deduced that if Assumption 2 holds, there exists $\gamma_{u_{0}}>0$ such that for all $\gamma_{u}\geq\gamma_{u_{0}}$, $\mathcal{A}_{u}$ is Hurwitz stable. Since $K_{d,u}$, $R_{d,u}$, and $Q_{u}$ are bounded, it follows from Lemma 3 that $\Sigma_{u}(t)$ will converge to $\bar{\Sigma}_{u}$.
Item 2: In the scenario where $F_{d,u}\neq 0$, the convergence of $S(t)$ affects that of $\Sigma_{e}(t)$. If $A$ is a stable matrix, it can be concluded that $X(t)$, $S(t)$, and $\Sigma_{e}(t)$ converge based on Lemma 3. In the other case, if $F_{d,u}=0$, the convergence of $\Sigma_{e}(t)$ can be derived based on the boundedness of $Q$ and $R_{d}$, along with the principles of the Lyapunov equation theory. ∎
For the distributed filter with the nominal model, it is desired that the steady-state estimation error covariance $\bar{\Sigma}_{e}$ can be evaluated by the nominal performance index $\bar{\Sigma}_{u}$. Consequently, the bounds of the steady-state estimation error covariance are given in the following. Since $\mathcal{A}_{u}$ incorporates $\gamma_{u}$, and $\gamma_{u}$ may tend to infinity. Hence, it is crucial to identify the effect of $\gamma_{u}$.
Theorem 1.
Consider the system (1) and the nominal distributed filter (5). Suppose Assumptions 1-3 hold. Define $s=\|I_{nN}\otimes\mathcal{A}_{u}+A_{d,u}\otimes I_{nN}\|_{F}-\sqrt{N}\|\Delta A_{d}\|_{F}$ and $\chi=\|I_{nN}\otimes A_{d,u}+A_{d,u}\otimes I_{nN}\|_{F}-2\sqrt{nN}\|\Delta A_{d}\|_{F}$. If $\gamma_{u}\geq\gamma_{u_{0}}$, $s\neq 0$, and $\chi\neq 0$, then
  1. 1.
    $\text{Tr}(\bar{\Sigma}_{e})$ is bounded as follows
    $\max\{0,\text{Tr}(\bar{\Sigma}_{u})-\rho\}\leq\text{Tr}(\bar{\Sigma}_{e})\leq\text{Tr}(\bar{\Sigma}_{u})+\rho,$
    where
    $~{}~{}\rho$ $=\|(\text{vec}(I_{nN}))^{T}\mathcal{\bar{A}}^{-1}_{u}(K_{d,u}\otimes K_{d,u})\|_{2}$ (23)
    ×$~{}~{}~{}~{}\times\sqrt{\sum^{N}_{j=1}\|\Delta R_{j}\|^{2}_{F}}+\|(\text{vec}(I_{nN}))^{T}\mathcal{\bar{A}}^{-1}_{u}\|_{2}$
    ×$~{}~{}~{}~{}\times\Big{(}N\sqrt{\|\Delta Q\|^{2}_{F}}+2\|F_{d,u}\|_{F}\|\bar{S}\|_{F}\Big{)},$
    $~{}~{}~{}~{}\|\bar{S}\|_{F}$ (24)
    $\leq\frac{\sqrt{nN}\|F_{d,u}\|_{F}\|\bar{X}\|_{F}+N\|Q_{u}\|_{F}+N\|\Delta Q\|_{F}}{\text{abs}(s)},~{}~{}$
    and
    $\|\bar{X}\|_{F}\leq\frac{N\|Q_{u}\|_{F}+N\|\Delta Q\|_{F}}{\text{abs}(\chi)}.{}{}{}{}{}{}{}{}{}$ (25)
  2. 2.
    $\rho$ is asymptotic decay with respect to $\gamma_{u}$, where
    $\|(\text{vec}(I_{nN}))^{T}\mathcal{\bar{A}}^{-1}_{u}\|_{2}=\sqrt{a_{1}+\frac{b_{1}}{\gamma_{u}}+\frac{c_{1}}{\gamma^{2}_{u}}+O\big{(}\frac{1}{\gamma^{3}_{u}}\big{)}},{}{}{}{}{}{}{}{}$ (26)
    $~{}~{}~{}~{}\|(\text{vec}(I_{nN}))^{T}\mathcal{\bar{A}}^{-1}_{u}(K_{d,u}\otimes K_{d,u})\|_{2}$ (27)
    $=\sqrt{a_{2}+\frac{b_{2}}{\gamma_{u}}+\frac{c_{2}}{\gamma^{2}_{u}}+O\big{(}\frac{1}{\gamma^{3}_{u}}\big{)}},~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}$
    and $a_{1}$, $b_{1}$, $c_{1}$, $a_{2}$, $b_{2}$, and $c_{2}$ are constants, with $a_{1},b_{1},a_{2},b_{2}>0$.
Proof.
Item 1: By defining $\bar{D}=-F_{d,u}\bar{S}^{T}-\bar{S}F^{T}_{d,u}$, (22) can be rewritten as
$~{}~{}~{}~{}\mathcal{A}_{u}\bar{\Sigma}_{e}+\bar{\Sigma}_{e}\mathcal{A}^{T}_{u}+K_{d,u}R_{d,u}K^{T}_{d,u}+U_{N}\otimes Q_{u}$ (28)
$=K_{d,u}\Delta R_{d}K^{T}_{d,u}+U_{N}\otimes\Delta Q+\bar{D}.~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}$
Now, consider the vectorization of (28). According to the identity $\text{vec}(XYZ)=(Z^{T}\otimes X)\text{vec}(Y)$, one has
$~{}~{}~{}~{}\mathcal{\bar{A}}_{u}\text{vec}(\bar{\Sigma}_{e})+\text{vec}(K_{d,u}R_{d,u}K^{T}_{d,u}+U_{N}\otimes Q_{u})$ (29)
$=(K_{d,u}\otimes K_{d,u})\text{vec}(\Delta R_{d})+\text{vec}(U_{N}\otimes\Delta Q)+\text{vec}(\bar{D}),$
where $\mathcal{\bar{A}}_{u}=(I_{Nn}\otimes\mathcal{A}_{u})+(\mathcal{A}_{u}\otimes I_{Nn})$. Similarly to (29), (21) can be calculated as
$\mathcal{\bar{A}}_{u}\text{vec}(\bar{\Sigma}_{u})+\text{vec}(K_{d,u}R_{d,u}K^{T}_{d,u}+U_{N}\otimes Q_{u})=0.$ (30)
By combining (29) and (30), it follows
$\mathcal{\bar{A}}_{u}\text{vec}(\bar{\Sigma}_{e})$ $=\mathcal{\bar{A}}_{u}\text{vec}(\bar{\Sigma}_{u})+(K_{d,u}\otimes K_{d,u})\text{vec}(\Delta R_{d})~{}~{}~{}$
$~{}~{}~{}~{}+\text{vec}(U_{N}\otimes\Delta Q)+\text{vec}(\bar{D}).$
By utilizing the properties of the Kronecker sum and the fact that $\mathcal{A}_{u}$ is Hurwitz stable, it can be deduced that $\mathcal{\bar{A}}_{u}$ is also Hurwitz stable. Hence, it holds
$\text{vec}(\bar{\Sigma}_{e})$ $=\text{vec}(\bar{\Sigma}_{u})+\mathcal{\bar{A}}^{-1}_{u}(K_{d,u}\otimes K_{d,u})\text{vec}(\Delta R_{d})$ (31)
$~{}~{}~{}~{}+\mathcal{\bar{A}}^{-1}_{u}\text{vec}(U_{N}\otimes\Delta Q)+\mathcal{\bar{A}}^{-1}_{u}\text{vec}(\bar{D}).$
By utilizing the identity $\text{Tr}(X)=(\text{vec}(I))^{T}\text{vec}(X)$, (31) can be rewritten as
$\text{Tr}(\bar{\Sigma}_{e})$ $=\text{Tr}(\bar{\Sigma}_{u})+\rho_{b},$ (32)
where
$\rho_{b}$ $=(\text{vec}(I_{nN}))^{T}\mathcal{\bar{A}}^{-1}_{u}(K_{d,u}\otimes K_{d,u})\text{vec}(\Delta R_{d})$
$~{}~{}~{}~{}+(\text{vec}(I_{nN}))^{T}\mathcal{\bar{A}}^{-1}_{u}\text{vec}(U_{N}\otimes\Delta Q)$
$~{}~{}~{}~{}+(\text{vec}(I_{nN}))^{T}\mathcal{\bar{A}}^{-1}_{u}\text{vec}(\bar{D}).$
Then, based on the properties of (32), one has $\|\text{Tr}(\bar{\Sigma}_{e})-\text{Tr}(\bar{\Sigma}_{u})\|_{2}=\|\rho_{b}\|_{2}$. According to the identity $\|\text{vec}(X)\|_{F}=\|X\|_{F}$ and $\|X\|_{2}\leq\|X\|_{F}$, one has
$\|\rho_{b}\|_{2}$ $\leq\|(\text{vec}(I_{nN}))^{T}\mathcal{\bar{A}}^{-1}_{u}(K_{d,u}\otimes K_{d,u})\|_{2}$ (33)
×$~{}~{}~{}~{}\times\sqrt{\sum^{N}_{j=1}\|\Delta R_{j}\|^{2}_{F}}+\|(\text{vec}(I_{nN}))^{T}\mathcal{\bar{A}}^{-1}_{u}\|_{2}$
×$~{}~{}~{}~{}\times\Big{(}N\sqrt{\|\Delta Q\|^{2}_{F}}+\|\bar{D}\|_{F}\Big{)}.$
By defining $\rho$ as stated in (23), it can be deduced that $\|\text{Tr}(\bar{\Sigma}_{e})-\text{Tr}(\bar{\Sigma}_{u})\|_{2}\leq\rho$.
Next, we derive the upper bound of $\|\bar{D}\|_{F}$. Based on $\bar{D}=-F_{d,u}\bar{S}^{T}-\bar{S}F^{T}_{d,u}$, we have
$\|\bar{D}\|_{F}\leq 2\|F_{d,u}\|_{F}\|\bar{S}\|_{F}.$ (34)
According to the expression of $\bar{S}$ below (22) and applying the vectorization operation to it, it follows that
$~{}~{}~{}~{}(I_{nN}\otimes\mathcal{A}_{u}+A_{d,u}\otimes I_{nN})\text{vec}(\bar{S})$ (35)
$=(\Delta A_{d}\otimes I_{nN})\text{vec}(\bar{S})-(I_{nN}\otimes F_{d,u})\text{vec}(\bar{X})$
$~{}~{}~{}~{}-\text{vec}(U_{N}\otimes Q).$
By taking the Frobenius norm of (35), if $\|I_{nN}\otimes\mathcal{A}_{u}+A_{d,u}\otimes I_{nN}\|_{F}-\sqrt{nN}\|\Delta A_{d}\|_{F}>0$, then
$~{}~{}~{}~{}\|I_{nN}\otimes\mathcal{A}_{u}+A_{d,u}\otimes I_{nN}\|_{F}\|\bar{S}\|_{F}$ (36)
$\leq\sqrt{nN}\|\Delta A_{d}\|_{F}\|\bar{S}\|_{F}+\sqrt{nN}\|F_{d,u}\|_{F}\|\bar{X}\|_{F}$
$~{}~{}~{}~{}+\|U_{N}\otimes Q\|_{F}.$
If $\|I_{nN}\otimes\mathcal{A}_{u}+A_{d,u}\otimes I_{nN}\|_{F}-\sqrt{nN}\|\Delta A_{d}\|_{F}<0$, then
$\sqrt{nN}\|\Delta A_{d}\|_{F}\|\bar{S}\|_{F}\leq\|I_{nN}\otimes\mathcal{A}_{u}+A_{d,u}\otimes I_{nN}\|_{F}\|\bar{S}\|_{F}$ (37)
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}+\sqrt{nN}\|F_{d,u}\|_{F}\|\bar{X}\|_{F}+\|U_{N}\otimes Q\|_{F}.$
By combining (36) and (37), it can be concluded that if $\|I_{nN}\otimes\mathcal{A}_{u}+A_{d,u}\otimes I_{nN}\|_{F}-\sqrt{nN}\|\Delta A_{d}\|_{F}\neq 0$, then $\|\bar{S}\|_{F}$ can be bounded as
$~{}~{}~{}~{}\|\bar{S}\|_{F}$ (38)
$\leq\frac{\sqrt{nN}\|F_{d,u}\|_{F}\|\bar{X}\|_{F}+N\|Q_{u}\|_{F}+N\|\Delta Q\|_{F}}{\text{abs}(\|I_{nN}\otimes\mathcal{A}_{u}+A_{d,u}\otimes I_{nN}\|_{F}-\sqrt{nN}\|\Delta A_{d}\|_{F})}.$
Similarly, based on $A_{d}\bar{X}+\bar{X}A^{T}_{d}+U_{N}\otimes Q=0$, we can obtain that if $\|I_{nN}\otimes A_{d,u}+A_{d,u}\otimes I_{nN}\|_{F}-2\sqrt{nN}\|\Delta A_{d}\|_{F}\neq 0$, then
$~{}~{}~{}~{}\|\bar{X}\|_{F}$ (39)
$\leq\frac{N\|Q_{u}\|_{F}+N\|\Delta Q\|_{F}}{\text{abs}(\|I_{nN}\otimes A_{d,u}+A_{d,u}\otimes I_{nN}\|_{F}-2\sqrt{nN}\|\Delta A_{d}\|_{F})}.$
Based on (33), (34), (38), and (39), the results can be proven.
Item 2: First, we transform the problem of computing $(\text{vec}(I_{nN}))^{T}\mathcal{\bar{A}}^{-1}_{u}$ into finding the solution $X$ of the Lyapunov equation $\mathcal{A}^{T}_{u}X+X\mathcal{A}_{u}=I_{nN}$. Since $(\text{vec}(I_{nN}))^{T}\mathcal{\bar{A}}^{-1}_{u}=(\mathcal{\bar{A}}^{-T}_{u}\text{vec}(I_{nN}))^{T}$, our focus is on $\mathcal{\bar{A}}^{-T}_{u}\text{vec}(I_{nN})$. Given that $\lambda(\mathcal{A}_{u})\cap\lambda(-\mathcal{A}^{T}_{u})=\emptyset$, $X$ is the unique solution of the Lyapunov equation $\mathcal{A}^{T}_{u}X+X\mathcal{A}_{u}=I_{nN}$. Thus, we have $\text{vec}(X)=\mathcal{\bar{A}}^{-T}_{u}\text{vec}(I_{nN})$, where $\mathcal{\bar{A}}^{T}_{u}=(I_{nN}\otimes\mathcal{A}^{T}_{u})+(\mathcal{A}^{T}_{u}\otimes I_{nN})$.
Second, we derive an analytic solution for $X$ in terms of $\gamma_{u}$. Based on the property of the communication topology $\mathcal{L}$, there exist a transformation matrix $\bar{T}$ such that $\tilde{\Lambda}=\bar{T}\mathcal{L}\bar{T}^{T}=\text{diag}\{0,\lambda_{2},\ldots,\lambda_{N}\}$, where $\lambda_{2},\ldots,\lambda_{N}$ are the positive eigenvalues of $\mathcal{L}$, and the first row of $\bar{T}$ is the vector with all elements equal to $1$. Denote $T=\bar{T}\otimes I_{n}$, then $\gamma_{u}T(\mathcal{L}\otimes P_{u}(\infty))T^{T}=\gamma_{u}(\tilde{\Lambda}\otimes P_{u}(\infty))$. We can rewrite the Lyapunov equation as
$(\tilde{G}^{T}-\gamma_{u}(\tilde{\Lambda}\otimes P_{u}(\infty)))\tilde{X}$ (40)
$+\tilde{X}(\tilde{G}-\gamma_{u}(\tilde{\Lambda}\otimes P_{u}(\infty)))=I_{nN},$
where $\tilde{G}=TG_{d,u}T^{T}$ and $\tilde{X}=TXT^{T}$. Split the matrix $\tilde{G}$ into blocks as follows
$\tilde{G}=\left[\begin{array}[]{cc}\tilde{G}_{11}&\tilde{G}_{12}\\ \tilde{G}^{T}_{12}&\tilde{G}_{22}\end{array}\right],$ (41)
where $\tilde{G}_{11}\in\mathcal{R}^{n\times n}$, $\tilde{G}_{12}\in\mathcal{R}^{n\times n(N-1)}$, and $\tilde{G}_{22}\in\mathcal{R}^{n(N-1)\times n(N-1)}$. Denote $\Lambda=\text{diag}\{\lambda_{2},\ldots,\lambda_{N}\}$, and we have
$(\tilde{G}-\gamma_{u}(\tilde{\Lambda}\otimes P_{u}(\infty)))=\left[\begin{array}[]{cc}\tilde{G}_{11}&\tilde{G}_{12}\\ \tilde{G}^{T}_{12}&\tilde{G}_{22}-\gamma_{u}\Lambda\otimes P_{u}(\infty)\end{array}\right].$ (42)
We can define the matrix $\tilde{X}$ as a block matrix with the same dimensions as the sub-blocks of $\tilde{G}$ in (41). By examining these block matrices, it can be hypothesized that the solution of $\tilde{X}$ is analytic in $\gamma_{u}$. The analytic form can be determined by utilizing the block matrices in (40) and analyzing four equalities. The solution of $\tilde{X}(\gamma_{u})$ can be given as
$\tilde{X}(\gamma_{u})=\left[\begin{array}[]{cc}\tilde{X}_{11,1}+\frac{1}{\gamma_{u}}\tilde{X}_{11,2}+O(\frac{1}{\gamma^{2}_{u}})&\frac{1}{\gamma_{u}}\tilde{X}_{12}+O(\frac{1}{\gamma^{2}_{u}})\\ \frac{1}{\gamma_{u}}\tilde{X}^{T}_{12}+O(\frac{1}{\gamma^{2}_{u}})&\frac{1}{\gamma_{u}}\tilde{X}_{22}+O(\frac{1}{\gamma^{2}_{u}})\end{array}\right],$ (43)
where $\tilde{X}_{22}$ is the unique solution of $(\Lambda\otimes P_{u}(\infty))\tilde{X}_{22}+\tilde{X}_{22}(\Lambda\otimes P_{u}(\infty))=I_{n(N-1)}$, $\tilde{X}_{11,1}$ is the unique solution of $\tilde{G}_{11}\tilde{X}_{11,1}+\tilde{X}_{11,1}\tilde{G}_{11}=I_{n}$, $\tilde{X}_{12}$ is $\tilde{X}_{12}=\tilde{X}_{11,1}\tilde{G}_{12}(\Lambda\otimes P_{u}(\infty))^{-1}$, and $\tilde{X}_{11,2}$ is the unique solution of $\tilde{G}_{11}\tilde{X}_{11,2}+\tilde{X}_{11,2}\tilde{G}_{11}+\tilde{G}_{12}\tilde{X}^{T}_{12}+\tilde{X}_{12}\tilde{G}^{T}_{12}=0$.
Finally, since $X(\gamma_{u})=T^{T}\tilde{X}(\gamma_{u})T$, multiplying a constant matrix into $\tilde{X}(\gamma_{u})$ does not influence the structure of each term of $X(\gamma_{u})$ in terms of $\gamma_{u}$. We can conclude
$\|\text{vec}(X(\gamma_{u}))\|_{2}=\sqrt{a_{1}+\frac{b_{1}}{\gamma_{u}}+\frac{c_{1}}{\gamma^{2}_{u}}+O\big{(}\frac{1}{\gamma^{3}_{u}}\big{)}},$ (44)
where $a_{1}$, $b_{1}$, and $c_{1}$ are constants with $a_{1},b_{1}>0$. Similarly, $\|(\text{vec}(I_{nN}))^{T}\mathcal{\bar{A}}^{-1}_{u}(K_{d,u}\otimes K_{d,u})\|_{2}$ has a similar form as (44).
Remark 3.
The inequality condition in Theorem 1 can be easily satisfied by slightly increasing the deviation $\|\Delta A_{d}\|_{F}$. Additionally, $\bar{X}$ can be evaluated and replaced by using its known upper bound. Theorem 1 provides bounds of the steady-state performance of the distributed continuous-time filter. These bounds can be evaluated utilizing the nominal performance index and the knowledge of the Frobenius norms of $\Delta A$, $\Delta C_{i}$, $\Delta Q$, and $\Delta R_{i}$. Item 2 reveals the effect of the consensus parameter $\gamma_{u}$ on the upper bound of the estimation error covariance, and provides the asymptotic decay behavior of this bound in terms of $\gamma_{u}$.
Furthermore, it is desired to establish the lower bound of $\text{Tr}(\bar{\Sigma}_{u})$, which can also be utilized to evaluate the estimation error covariance. The lower bound of $\text{Tr}(\bar{\Sigma}_{u})$ is provided as follows.
Proposition 4.
Consider the system (1) and the nominal distributed filter (5). Suppose Assumptions 1-3 hold. If $\gamma_{u}\geq\gamma_{u_{0}}$, then the lower bound of $\text{Tr}(\bar{\Sigma}_{u})$ is
$\text{Tr}(\bar{\Sigma}_{u})$ $\geq\frac{\text{Tr}(K_{d,u}R_{d,u}K^{T}_{d,u}+U_{N}\otimes Q_{u})}{2\text{Tr}(-\mathcal{A}_{u_{0}})+2(\gamma_{u}-\gamma_{u_{0}})\text{Tr}(\mathcal{L})\text{Tr}(P_{u}(\infty))}.$ (45)
Proof.
Taking the trace operation on (21), one has
$~{}~{}~{}~{}\text{Tr}(\mathcal{A}_{u}\bar{\Sigma}_{u})+\text{Tr}(\bar{\Sigma}_{u}\mathcal{A}^{T}_{u})$ (46)
$=-\text{Tr}(K_{d,u}R_{d,u}K^{T}_{d,u})-\text{Tr}(U_{N}\otimes Q_{u}).$
Based on the property of the trace $\text{Tr}(XY)=\text{Tr}(YX)$, it follows
$\text{Tr}(\mathcal{A}^{T}_{u})$ $=\text{Tr}(\bar{\Sigma}^{-1}_{u}\bar{\Sigma}_{u}\mathcal{A}^{T}_{u})$ (47)
$=\text{Tr}(\bar{\Sigma}_{u}\mathcal{A}^{T}_{u}\bar{\Sigma}^{-1}_{u})$
$=\text{Tr}(\bar{\Sigma}^{-1}_{u}\mathcal{A}_{u}\bar{\Sigma}_{u}).$
According to (47) and (21), it holds
$2\text{Tr}(\mathcal{A}^{T}_{u})$ $=\text{Tr}(\bar{\Sigma}^{-1}_{u}(\bar{\Sigma}_{u}\mathcal{A}^{T}_{u}+\mathcal{A}_{u}\bar{\Sigma}_{u}))$ (48)
$=-\text{Tr}(\bar{\Sigma}^{-1}_{u}(K_{d,u}R_{d,u}K^{T}_{d,u}+U_{N}\otimes Q_{u})).$
By utilizing Lemma 2 and the fact that $K_{d,u}R_{d,u}K^{T}_{d,u}+U_{N}\otimes Q_{u}>0$ and $\bar{\Sigma}_{u}>0$, one has
$~{}~{}~{}~{}\text{Tr}(\bar{\Sigma}^{-1}_{u}(K_{d,u}R_{d,u}K^{T}_{d,u}+U_{N}\otimes Q_{u}))$ (49)
$\geq(\text{Tr}(\bar{\Sigma}_{u}))^{-1}\text{Tr}(K_{d,u}R_{d,u}K^{T}_{d,u}+U_{N}\otimes Q_{u}).~{}~{}~{}~{}~{}$
By combining (48) and (49), it follows
$~{}~{}2\text{Tr}(-\mathcal{A}_{u})\text{Tr}(\bar{\Sigma}_{u})$ $\geq\text{Tr}(K_{d,u}R_{d,u}K^{T}_{d,u}+U_{N}\otimes Q_{u}).$ (50)
Since all eigenvalues of $\mathcal{A}_{u}$ are negative, (50) can be rewritten as
$\text{Tr}(\bar{\Sigma}_{u})$ $\geq\frac{1}{2}(\text{Tr}(-\mathcal{A}_{u}))^{-1}\text{Tr}(K_{d,u}R_{d,u}K^{T}_{d,u}+U_{N}\otimes Q_{u}).$ (51)
Note that there exists
$-\mathcal{A}_{u}=-\mathcal{A}_{u_{0}}+(\gamma_{u}-\gamma_{u_{0}})(\mathcal{L}\otimes P_{u}(\infty)).$ (52)
Based on (52), it holds $\text{Tr}(-\mathcal{A}_{u})=\text{Tr}(-\mathcal{A}_{u_{0}})+(\gamma_{u}-\gamma_{u_{0}})\text{Tr}(\mathcal{L})\text{Tr}(P_{u}(\infty))$. Therefore, (51) can be simplified as
$\text{Tr}(\bar{\Sigma}_{u})$ $\geq\frac{\text{Tr}(K_{d,u}R_{d,u}K^{T}_{d,u}+U_{N}\otimes Q_{u})}{2\text{Tr}(-\mathcal{A}_{u_{0}})+2(\gamma_{u}-\gamma_{u_{0}})\text{Tr}(\mathcal{L})\text{Tr}(P_{u}(\infty))}.$ (53)
Remark 4.
By combining Proposition 4 and Theorem 1, the lower bound of $\text{Tr}(\bar{\Sigma}_{e})$ can be directly evaluated without the need for computing $\text{Tr}(\bar{\Sigma}_{u})$. The relations between the lower bound of $\text{Tr}(\bar{\Sigma}_{u})$ and the model parameters are presented in Proposition 4, and the consensus parameter is separated in the lower bound. It is shown that as the consensus parameter tends to infinity, the lower bound converges to 0.
3.2 Divergence Analysis
This subsection considers the divergence of the distributed continuous-time filter due to the incorrect process noise covariance.
Theorem 2.
Consider the system (1) and the nominal distributed filter (5). Suppose Assumption 1 hold. Let $r$ be a real number, $j$ be the imaginary unit, and $e$ be a vector with the appropriate dimension. If $A^{T}_{u}e=rje$ and $Q_{u}e=0$, then
  1. 1.
    $rj$ is an eigenvalue of $\mathcal{A}^{T}_{u}$ with the eigenvector $1_{N}\otimes e$, i.e., $\mathcal{A}^{T}_{u}(1_{N}\otimes e)=rj(1_{N}\otimes e)$, for any consensus parameter $\gamma_{u}$ and any undirected connected topology $\mathcal{L}$.
  2. 2.
    If $\Delta F_{d,u}=0$ and $Qe\neq 0$, it holds $\lim_{t\to\infty}(1^{T}_{N}\otimes e^{*})\Sigma_{e}(t)(1_{N}\otimes e)\to\infty.$
Proof.
Item 1: If $A^{T}_{u}e=rje$ and $Q_{u}e=0$, one has $e^{*}A_{u}=-rje^{*}$. By pre-multiplying with $e^{*}$ and post-multiplying with $e$ in (6), the subsequent expression is derived
$e^{*}A_{u}P_{u}(\infty)e+e^{*}P_{u}(\infty)A^{T}_{u}e+e^{*}Q_{u}e$ (54)
$-e^{*}P_{u}(\infty)C^{T}_{c,u}R^{-1}_{d,u}C_{c,u}P_{u}(\infty)e=0.$
Based on $A^{T}_{u}e=rje$, $e^{*}A_{u}=-rje^{*}$, and $Q_{u}e=0$, (54) can be simplified as
$~{}~{}~{}~{}e^{*}P_{u}(\infty)C^{T}_{c,u}R^{-1}_{d,u}C_{c,u}P_{u}(\infty)e$ (55)
$=\sum^{N}_{j=1}e^{*}P_{u}(\infty)C^{T}_{j,u}R^{-1}_{j,u}C_{j,u}P_{u}(\infty)e~{}~{}~{}~{}~{}$
$=0.$
Since $e^{*}P_{u}(\infty)C^{T}_{j,u}R^{-1}_{j,u}C_{j,u}P_{u}(\infty)e\geq 0$, it can be concluded that
$R^{-1}_{i,u}C_{i,u}P_{u}(\infty)e=0,i=1,\ldots,N.$ (56)
Utilizing (7) and the property of the undirected connected graph, post-multiply $\mathcal{A}^{T}_{u}$ by the vector $1_{N}\otimes e$, yielding
$\mathcal{A}^{T}_{u}(1_{N}\otimes e)$ $=G^{T}_{d,u}(1_{N}\otimes e)-\gamma_{u}(\mathcal{L}^{T}\otimes P_{u}(\infty))(1_{N}\otimes e)$ (57)
$=[\bar{e}^{T}_{1},\ldots,\bar{e}^{T}_{N}]^{T},$
where $\bar{e}_{i}=A^{T}_{u}e-NC^{T}_{i,u}R^{-1}_{i,u}C_{i,u}P_{u}(\infty)e$. Exploiting the fact $A^{T}_{u}e=rje$ and (56), it is evident that $\mathcal{A}^{T}_{u}(1_{N}\otimes e)=rj(1_{N}\otimes e)$. Consequently, $rj$ is a purely imaginary eigenvalue of $\mathcal{A}_{u}$.
Item 2: Item 1 and (56) lead to the conclusion that $R^{-1}_{i,u}C_{i,u}P_{u}(\infty)e=0$. Considering the expression $K_{d,u}=\text{diag}\{K_{1,u},\ldots,K_{N,u}\}$, it can be inferred that $K^{T}_{d,u}(1_{N}\otimes e)=0$. Additionally, Item 1 demonstrates that $\mathcal{A}^{T}_{u}(1_{N}\otimes e)=rj(1_{N}\otimes e)$. By pre-multiplying with $1^{T}_{N}\otimes e^{*}$, and post-multiplying with $1_{N}\otimes e$ in (18), one has
$~{}~{}~{}~{}(1^{T}_{N}\otimes e^{*})\dot{\Sigma}_{e}(t)(1_{N}\otimes e)$ (58)
$=(1^{T}_{N}\otimes e^{*})\mathcal{A}_{u}\Sigma_{e}(t)(1_{N}\otimes e)$
$~{}~{}~{}~{}+(1^{T}_{N}\otimes e^{*})\Sigma_{e}(t)\mathcal{A}^{T}_{u}(1_{N}\otimes e)$
$~{}~{}~{}~{}+(1^{T}_{N}\otimes e^{*})K_{d,u}R_{d}K^{T}_{d,u}(1_{N}\otimes e)$
$~{}~{}~{}~{}+(1^{T}_{N}\otimes e^{*})U_{N}\otimes Q(1_{N}\otimes e)$
$\geq(1^{T}_{N}\otimes e^{*})U_{N}\otimes Q(1_{N}\otimes e)>0.$
It can be observed that (58) as $t$ tends to infinity, $(1^{T}_{N}\otimes e^{*})\Sigma_{e}(t)(1_{N}\otimes e)$ will continuously increase. Consequently, this term diverges, and the proof is concluded. ∎
Remark 5.
Item 1 highlights that if $A^{T}_{u}$ possesses an eigenvalue $rj$ on the imaginary axis, which is linked to the right eigenvector $e$ in the right null space of $Q_{u}$, then $\mathcal{A}^{T}_{u}$ will also have $rj$ as its eigenvalue with the right eigenvector $1_{N}\otimes e$. Item 2 reveals that an incorrect process noise covariance can lead to the divergence of the estimation error covariance in the distributed continuous-time filter, even though the nominal performance index may converge. Therefore, it is advisable to ensure that the nominal model can be designed in such a way that Assumption 2 holds, thus avoiding the presence of pure imaginary eigenvalues. The effect of the consensus parameter on our conclusion is eliminated. It is important to note that the prerequisite for the communication graph to be undirected must be fulfilled. Otherwise, this inference may not be valid when assuming a directed graph.
Remark 6.
When the communication topology is mismatched, it can lead to the divergence of the estimation error covariance. This may violate the condition presented in Proposition 2, resulting in the instability of $\mathcal{A}_{u}$. When the mismatched topology is denoted as $\mathcal{L}_{u}$, we can use the lower bound of $\lambda_{N-1}(\mathcal{L}_{u})$ to compute $\gamma_{u_{0}}$ in (19) to guarantee the filter performance.
3.3 Relations between $\Sigma_{u}(t)$ and $\Sigma_{e}(t)$
Section 3.1 shows that the bounds of $\text{Tr}(\bar{\Sigma}_{e})$ can be evaluated by $\text{Tr}(\bar{\Sigma}_{u})$ and the Frobenius norms of the parameter deviation information. Furthermore, this subsection explores the relation between $\Sigma_{u}(t)$ and $\Sigma_{e}(t)$ by relying on the information about the magnitude relation between $R_{i,u}$ and $R_{i}$ as well as between $Q_{u}$ and $Q$. It is assumed that $\Delta A=0$ and $\Delta C_{i,u}=0,i=1,\ldots,N$.
First, define the difference $E_{ue}(t)$ between the nominal performance index $\Sigma_{u}(t)$ and the estimation error covariance $\Sigma_{e}(t)$ as $E_{ue}(t)=\Sigma_{u}(t)-\Sigma_{e}(t).$ Then, by utilizing (9) and (18), the differential equation of $\dot{E}_{ue}(t)$ can be computed as $\dot{E}_{ue}(t)=\mathcal{A}_{u}E_{ue}(t)+E_{ue}(t)\mathcal{A}^{T}_{u}+K_{d,u}\Delta R_{d}K^{T}_{d,u}+U_{N}\otimes\Delta Q.$ Define the time index $h$ satisfying $t\geq h\geq 0$. Drawing upon the principles of the linear system theory, a precise analytical solution can be calculated as
$E_{ue}(t)$ $=e^{\mathcal{A}_{u}(t-h)}E_{ue}(h)e^{\mathcal{A}^{T}_{u}(t-h)}+\int_{h}^{t}e^{\mathcal{A}_{u}(t-\tau)}$ (59)
×$~{}~{}~{}~{}\times(K_{d,u}\Delta R_{d}K^{T}_{d,u}+U_{N}\otimes\Delta Q)e^{\mathcal{A}^{T}_{u}(t-\tau)}d\tau.$
For further analysis, define
$\Delta D=K_{d,u}\Delta R_{d}K^{T}_{d,u}+U_{N}\otimes\Delta Q.$ (60)
The subsequent theorem shows the relations between $\Sigma_{u}(t)$ and $\Sigma_{e}(t)$. Moreover, it demonstrates how the upper bound of the spectral norm of the difference between $\Sigma_{u}(t)$ and $\Sigma_{e}(t)$ changes as time progresses.
Theorem 3.
Consider the system (1) and the nominal distributed filter (5). Suppose Assumptions 1-2 hold. If $\gamma_{u}\geq\gamma_{u_{0}}$ (as defined in (20)), $\Delta A=0$, and $\Delta C_{i,u}=0,i=1,\ldots,N$, it holds
  1. 1.
    If $\Delta D\geq 0$ and $E_{ue}(h)\geq 0$, then $\Sigma_{u}(t)\geq\Sigma_{e}(t)$. If $\Delta D\leq 0$ and $E_{ue}(h)\leq 0$, then $\Sigma_{u}(t)\leq\Sigma_{e}(t)$.
  2. 2.
    Define $\bar{\mu}(\mathcal{A}_{u_{0}})=\mu(\mathcal{A}_{u_{0}})+\mu(\mathcal{A}^{T}_{u_{0}})$ and $\mathcal{L}_{\gamma}=-(\gamma_{u}-\gamma_{u_{0}})(\mathcal{L}\otimes P_{u}(\infty))$. Then, $\|E_{ue}(t)\|_{2}$ has the following bound
    $\|E_{ue}(t)\|_{2}$ $\leq\|E_{ue}(h)\|_{2}e^{\bar{\mu}(\mathcal{A}_{u_{0}})(t-h)}e^{2\mu(\mathcal{L}_{\gamma})(t-h)}$ (61)
[…]
    Specifically, when the logarithmic norm is chosen as $\mu_{2}(\cdot)$, one has $\mu_{2}(\mathcal{L}_{\gamma})=0$.
Proof.
Item 1): By combining (59), these conclusions can be easily drawn.
Item 2): Taking the spectral norm of (59), it follows
$\|E_{ue}(t)\|_{2}$ $\leq\|E_{ue}(h)\|_{2}\|e^{\mathcal{A}_{u}(t-h)}\|_{2}\|e^{\mathcal{A}^{T}_{u}(t-h)}\|_{2}$ (62)
[…]
Based on Lemma 1, one has
$\|e^{\mathcal{A}_{u}(t-\tau)}\|_{2}\|e^{\mathcal{A}^{T}_{u}(t-\tau)}\|_{2}$ $\leq e^{\mu(\mathcal{A}_{u})(t-\tau)}e^{\mu(\mathcal{A}^{T}_{u})(t-\tau)}$ (63)
$\leq e^{(\mu(\mathcal{A}_{u})+\mu(\mathcal{A}^{T}_{u}))(t-\tau)}.$
Similarly to Theorem 1, based on the definition that $\mathcal{A}_{u_{0}}=G_{d,u}-\gamma_{u_{0}}(\mathcal{L}\otimes P_{u}(\infty))$, $\mathcal{L}_{\gamma}=-(\gamma_{u}-\gamma_{u_{0}})(\mathcal{L}\otimes P_{u}(\infty))$, and $\mathcal{A}_{u}=\mathcal{A}_{u_{0}}+\mathcal{L}_{\gamma}$, $e^{(\mu(\mathcal{A}_{u})+\mu(\mathcal{A}^{T}_{u}))(t-\tau)}$ can be rewritten as
$e^{(\mu(\mathcal{A}_{u})+\mu(\mathcal{A}^{T}_{u}))(t-\tau)}=e^{(\mu(\mathcal{A}_{u_{0}})+\mu(\mathcal{A}^{T}_{u_{0}}))(t-\tau)}e^{2\mu(\mathcal{L}_{\gamma})(t-\tau)}.$ (64)
By combining (63) and (64), (62) can be rewritten as
$\|E_{ue}(t)\|_{2}$ $\leq\|E_{ue}(h)\|_{2}e^{(\mu(\mathcal{A}_{u_{0}})+\mu(\mathcal{A}^{T}_{u_{0}}))(t-h)}e^{2\mu(\mathcal{L}_{\gamma})(t-h)}$ (65)
$~{}~{}~{}~{}+\|\Delta D\|_{2}\int_{h}^{t}e^{(\mu(\mathcal{A}_{u_{0}})+\mu(\mathcal{A}^{T}_{u_{0}}))(t-\tau)}$
×$~{}~{}~{}~{}\times e^{2\mu(\mathcal{L}_{\gamma})(t-\tau)}d\tau.$
Define $\bar{\mu}(\mathcal{A}_{u_{0}})=\mu(\mathcal{A}_{u_{0}})+\mu(\mathcal{A}^{T}_{u_{0}})$, and (61) can be obtained. Based on Assumption 1, the Laplacian matrix $\mathcal{L}$ is positive semi-definite with one zero eigenvalue. According to the definition of $\mu_{2}(\cdot)$ and the property of $\mathcal{L}_{\gamma}$, it can be deduced that $\mu_{2}(\mathcal{L}_{\gamma})=0$.
Remark 7.
On the one hand, Theorem 3 shows that the relative relation between $\Sigma_{u}(t)$ and $\Sigma_{e}(t)$ can be determined by utilizing $\Delta D$ and $E_{ue}(h)$. On the other hand, the evolution process of the upper bound of the spectral norm of the difference is presented. Item 2 provides a distinct separation of the consensus parameter in the upper bound. The logarithmic norm $\mu_{2}(\mathcal{L}_{\gamma})=0$ means that the magnitude of the coefficient of the exponential term is influenced by the logarithmic norm of $\mathcal{A}_{u_{0}}$ associated with $\gamma_{u_{0}}$, and the consensus parameter $\gamma_{u}$ satisfying $\gamma_{u}\geq\gamma_{u_{0}}$ have no impact on the bound. If $\Delta D=0$, Item 2 show the convergence rate between the nominal performance index $\Sigma_{u}(t)$ and the estimation error covariance $\Sigma_{e}(t)$.
4 Simulations
In this section, numerical experiments are conducted on vehicle tracking to validate the effectiveness of the theoretical results. The sensor network used in the experiments consists of six sensors, and its communication topology is illustrated in Fig. 1.
Figure 1: Illustration of the communication topology.
Consider that the vehicle is moving in a plane. Let $x(t)=[x^{T}_{1}(t),x^{T}_{2}(t),x^{T}_{3}(t),x^{T}_{4}(t)]^{T}$ denote the state, where $x_{1}(t)$ and $x_{3}(t)$ are the horizontal velocity and the vertical velocity, respectively, and $x_{2}(t)$ and $x_{4}(t)$ are the horizontal position and the vertical position, respectively. The vehicle dynamics is described by $A=[\bar{A}_{11},0;0,\bar{A}_{22}]$, where $\bar{A}_{11}=\bar{A}_{22}=[0,0;1,0]$, The process noise covariance is given as   $Q=\text{diag}\{0.03,0.03,0.03,0.03\}$. Two kinds of sensors are deployed in the sensor network, and their measurement matrices are defined as $C_{1}=[0,1,0,0]$, $C_{2}=[0,2,0,0]$, $C_{3}=[0,3,0,0]$, $C_{4}=[0,0,0,1]$, $C_{5}=[0,0,0,2]$, and $C_{6}=[0,0,0,3]$. The measurement noise covariances are set as $R_{i}=0.2,~{}i=1,\ldots,6$. The initial state is $x(0)=[0.2;1;0.2;1]$, and the initial covariance is $\Sigma_{i,e}(0)=\text{diag}\{0.1,0.1,0.1,0.1\}$. The Monte Carlo method is adopted to evaluate the performance of the distributed filter, and the mean square error (MSE) is utilized as a metric, given by $\text{MSE}(t)=\frac{1}{MN}\sum^{N}_{i=1}\sum^{M}_{l=1}\|\hat{x}^{(l)}_{i,u}(t)-x^{(l)}(t)\|^{2}_{2},$ where $N$ is the sensor number and $M$ is the trail number.
To demonstrate the theoretical results, three cases are considered. Case 1 validates the theoretical theories of the bounds of the estimation error covariance, the bounds of the nominal performance index, the estimation error covariance, and the collective observability condition. Case 2 illustrates the divergence of the distributed filter caused by an incorrect process noise covariance. Case 3 shows the relations between $\Sigma_{u}(t)$ and $\Sigma_{e}(t)$.
Case 1: Consider the bounds of the estimation error covariance. The vehicle dynamics is described by $A=[\bar{A}_{11},0;0,\bar{A}_{22}]$, where $\bar{A}_{11}=\bar{A}_{22}=[-0.1,0;1,-0.1]$, $A_{u}=A+0.1I_{4}$, $C_{1,u}=[0,1.1,0,0]$, $C_{2,u}=[0,2.1,0,0]$, $C_{3,u}=[0,3.1,0,0]$, $C_{4,u}=[0,0,0,1.1]$, $C_{5,u}=[0,0,0,2.1]$, $C_{6,u}=[0,0,0,3.1]$, $Q_{u}=0.05I_{4}$, and $R_{i,u}=0.3,~{}i=1,\ldots,6$. Fig. 2 displays $\text{Tr}(\bar{\Sigma}_{e})$, the upper bound 1 of $\text{Tr}(\bar{\Sigma}_{e})$ ($\text{Tr}(\bar{\Sigma}_{u})+\rho(\gamma_{u})$) referring to (23), the upper bound 2 of $\text{Tr}(\bar{\Sigma}_{e})$ ($\text{Tr}(\bar{\Sigma}_{u})+\rho(\gamma_{u}-0.5)$) referring to (23) and (26), the lower bound of $\text{Tr}(\bar{\Sigma}_{u})$ referring to (45), and MSE with the increasing consensus parameter $\gamma_{u}$. It is shown that
  1. 1.
    The curves of MSE and $\text{Tr}(\bar{\Sigma}_{e})$ exhibit nearly overlapping, demonstrating the correctness of the theoretical value $\bar{\Sigma}_{e}$.
  2. 2.
    As the consensus parameter $\gamma_{u}$ increases, five indices decline. Moreover, both MSE and $\text{Tr}(\bar{\Sigma}_{e})$ are constrained by the upper bound 1 and the upper bound 2, and the asymptotic decay of the upper bound is demonstrated, as established Theorem 1.
Figure 2: MSE, $\text{Tr}(\bar{\Sigma}_{e})$, the upper bound 1 of $\text{Tr}(\bar{\Sigma}_{e})$, the upper bound 2 of $\text{Tr}(\bar{\Sigma}_{e})$, and the lower bound of $\text{Tr}(\bar{\Sigma}_{u})$ with the increasing consensus parameter in Case 1.
Figure 3: Illustration of the divergence caused by the incorrect covariance with the increasing time in Case 2.
Case 2: To verify the effectiveness of Theorem 2, the divergence of the distributed filter resulting from the incorrect covariance are considered. Let us establish the nominal process noise covariance as $Q_{u}=\text{diag}\{0,0.03,0.03,0.03\}$, and set the consensus parameter as $\gamma_{u}=10$. First, compute the eigenvalues and eigenvectors of $A_{u}$. All eigenvalues of $A^{T}_{u}$ are 00, with the corresponding eigenvectors are $[1,0,0,0]^{T}$, $[-1,0,0,0]^{T}$, $[0,0,1,0]^{T}$, and $[0,0,-1,0]^{T}$. Consider the condition $A^{T}_{u}e=rje$ and $Q_{u}e=0$. From this, it can be deduced that $r=0$, and $e=[1,0,0,0]^{T}$ or $e=[-1,0,0,0]^{T}$. Ultimately, it can be calculated that $\mathcal{A}^{T}_{u}(1_{N}\otimes e)=0$. Fig. 3 shows MSE of the distributed filter, and it can be observed that the actual estimation error has diverged due to utilizing an incorrect process noise covariance (See Theorem 2).
Case 3: Consider the relations between $\Sigma_{u}(t)$ and $\Sigma_{e}(t)$. The nominal parameters are set as $A_{u}=A$, $C_{i,u}=C_{i}$, $Q_{u}=\text{diag}\{0.1,0.1,0.1,0.1\}$, $R_{i,u}=0.3,~{}i=1,4$, and $R_{i,u}=0.2,~{}i=2,3,5,6$. The initial estimation error covariance and the initial performance index are chosen as the identity matrix. As stated in (60), it follows $\Delta D>0$.
Figure 4: Illustration of the relations between $\Sigma_{u}(t)$ and $\Sigma_{e}(t)$ with the increasing time in Case 3.
Fig. 4 exhibits the trace of the estimation error covariance and the nominal performance index. This figure illustrates that the nominal performance index provides an upper bound of the estimation error covariance, which aligns with the theoretical results (See Theorem 3). It means that the estimation error covariance of the distributed filter can be evaluated by utilizing the nominal performance index, if the nominal parameters are appropriately designed.
5 Conclusions
In this paper, the performance analysis of the distributed continuous-time filter in the presence of the modeling errors is conducted from two primary aspects. On the one hand, the convergence condition and the corresponding convergence analysis of two performance indices in the presence of the modeling errors are provided. Then, it is demonstrated that an incorrect noise covariance can lead to the divergence of the distributed filter. On the other hand, this paper focuses on the performance evaluation from the nominal performance index and the estimation error covariance. The bounds of the estimation error are derived by utilizing the nominal performance index and the model deviation information. The relative magnitude relations between them are also presented. These results shed light on the performance of the distributed filter in the presence of the modeling errors and provide guidance for engineering applications. In the future, we aim to develop the corresponding distributed algorithms to handle the effects of the modeling errors.
Appendix A Graph Theory
A sensor network’s nodes and communication links can be represented as the communication topology ${\mathcal{G}}(\mathcal{V},\mathcal{E})$, where the node set and the edge set are denoted as $\mathcal{V}=\{1,2,\ldots,N\}$ and $\mathcal{E}\subseteq\mathcal{V}\times\mathcal{V}$, respectively. For $i,j\in\mathcal{V}$, if the information can be transmitted from node $j$ to node $i$, node $j$ is called a neighbor of node $i$, denoted as $(j,i)$. The neighbor set of node $i$ is represented as $\mathcal{N}_{i}=\{j|(j,i)\in\mathcal{V}\}$, and $|\mathcal{N}_{i}|$ is the cardinality of the neighbors of node $i$. The adjacent matrix is $S=[s_{ij}]_{N\times N}$, where $s_{ij}=1$ if $(j,i)\in\mathcal{E}$, and $s_{ij}=0$ otherwise. The Laplacian matrix is defined as $\mathcal{L}=D-S$, where $D=\text{diag}\{|\mathcal{N}_{1}|,\ldots,|\mathcal{N}_{N}|\}$. If $(i,j)\in\mathcal{E}$ implies $(j,i)\in\mathcal{E}$, the edge $(i,j)$ is called undirected. If every edge is undirected, the communication graph is termed undirected. A directed path from node $i_{1}$ to node $i_{m}$ exists in the graph $G$, if there is a sequence of connected edges $(i_{k},i_{k+1}),k=1,\ldots,m-1$. The undirected communication graph is connected if there exists a path between every two nodes.
Appendix B Supporting Lemmas
Lemma 1.
[16] For matrices $A$ and $B$ with the appropriate dimensions, the following properties of the logarithmic norm $\mu(\cdot)$ hold:
  1. 1.
    $\mu(\gamma A)=\gamma\mu(A)$ for scalar $\gamma>0$,
  2. 2.
    $\mu(A+B)\leq\mu(A)+\mu(B)$,
  3. 3.
    $\|e^{tA}\|\leq e^{t\mu(A)}$ for $t\geq 0$.
Lemma 2.
[19] For matrices $A$ and $B$ with the appropriate dimensions, if $A>0$ and $B\geq 0$, then $\text{Tr}(AB)\leq\text{Tr}(A)\bar{\sigma}(B)\leq\text{Tr}(A)\text{Tr}(B),$ and $\text{Tr}(A^{-1}B)\geq(\bar{\sigma}(A))^{-1}\text{Tr}(B)\geq[\text{Tr}(A)]^{-1}\text{Tr}(B).$
Lemma 3.
[21] For matrices $A$, $B$, and $C$ with the appropriate dimensions, consider the Sylvester equation $AX+XB=C.$ There exists a unique solution $X$ if and only if $\lambda_{i}(A)+\lambda_{j}(B)\neq 0$ for $i=1,\ldots,m$ and $j=1,\ldots,n$.